# Gaussian Copula Models for Nonignorable Missing Data Using Auxiliary Marginal Quantiles

### v1

@@ 5 Analysis of North Carolina Lead Exposure Data
To illuminate the effect of the EHQL imputations further, we examine the empirical CDF of Blood_lead obtained by averaging across the $20$ completed datasets; this is displayed in Figure 7. Because it utilizes $\mathcal{A}$, the EHQL imputes values of missing Blood_lead measurements that are small relative to the observed values. Simultaneously, the percentiles of EoG math scores are higher for students missing Blood_lead than for students with recorded values. Consequently, imputing Blood_lead so that its completed-data distribution accords with $\mathcal{A}$ strengthens its negative association with Math_Score. We see similar strengthening of inverse associations with the other predictors except for RI, which has estimates mostly shrunk towards zero. RI is strongly correlated to NDI, which may help explain why its estimates attenuate. We note that a 95% credible interval for the copula correlation between Blood_lead and its missingness indicator is (-0.92,-0.91), offering additional evidence that the missingness in lead exposure measurements is MNAR.
@@ 6 Concluding Remarks
There are many topics worthy of future research. For example, the marginal quantiles may be known with uncertainty, for example, estimates from a probability sample. In cases where analysts desire a single inference, it may be possible to posit sampling distributions for the true quantiles that can be integrated into the model specification. Additionally, often data have survey weights. While there are methods for using auxiliary margins with survey-weighted data for categorical data models (Akande and Reiter, , 2022; Tang et al., , 2024), work is needed to develop methods for the EQL/EHQL copulas.
The North Carolina lead exposure analysis suggests that utilizing auxiliary marginal quantiles to handle nonignorable missing data can impact empirical findings. In particular, inferences drawn from the multiple imputations based on the EHQL copula suggest that lead exposure affects childhood cognitive development more adversely than might be concluded from a complete case analysis. More broadly, linked data like these are commonly used in public health studies and full of missing values. When the missingness may be MNAR, analysts can consider imputation strategies that leverage known marginal quantiles of the study variables to better inform health policy and intervention strategies.

### v2

@@ 5 Analysis of North Carolina Lead Exposure Data
Figure 7 displays the empirical CDF of Blood_lead obtained by averaging across the $20$ completed datasets. Because it utilizes $\mathcal{A}$, the EHQL imputes values of missing Blood_lead measurements that are small relative to the observed values. Simultaneously, the percentiles of EoG math scores are higher for students missing Blood_lead than for students with recorded values. Consequently, imputing Blood_lead so that its completed-data distribution accords with $\mathcal{A}$ strengthens its negative association with Math_Score. We see similar strengthening of inverse associations with the other predictors except for RI, which has estimates mostly shrunk towards zero. RI is strongly correlated to NDI, which may help explain why its estimates attenuate. We note that a 95% credible interval for the copula correlation between Blood_lead and its missingness indicator is (-0.92,-0.91), offering additional evidence that lead exposure measurements are MNAR.
@@ 6 Concluding Remarks
The marginal quantiles may be known with uncertainty, for example, estimates from a probability sample. When analysts desire a single inference, it may be possible to posit sampling distributions for the true quantiles that can be integrated into the model specification; this is a topic for future research. Additionally, often data have survey weights. While there are methods for using auxiliary margins with survey-weighted data for categorical data models (Akande and Reiter, 2022; Tang et al., 2024), work is needed to develop methods for the EQL/EHQL copulas.
The North Carolina lead exposure analysis suggests that utilizing auxiliary marginal quantiles to handle nonignorable missing data can impact empirical findings. In particular, inferences drawn from the multiple imputations based on the EHQL copula suggest that lead exposure affects childhood cognitive development more adversely than might be concluded from a complete case analysis. More broadly, linked data like these are commonly used in public health studies and full of missing values. When the missingness may be MNAR, analysts can consider imputation strategies that leverage known marginal quantiles of the study variables to better inform health policy.